# Slow Transition to Low-Dimensional Chaos in Heavy-Tailed Recurrent Neural Networks

### v1

@@ 4.3.1 Lyapunov spectrum shows compressed slow manifold in heavy-tailed RNNs
Second, we evaluate the participation ratio (PR), a widely used metric for gauging the effective dimensionality of neural population activity. It has been leveraged to quantify task-relevant low-dimensional subspaces and other properties of multi-unit neuronal recordings in behaving animals (Gao et al., 2017), and summarize the collective modes visited by recurrent spiking networks and reveal how these modes depend on local connectivity motifs (Recanatesi et al., 2019). PR measures how variance in population activity is distributed across the eigenmodes of the covariance matrix, providing a compact read-out of the number of degrees of freedom the network explores (Kramer and MacKinnon, 1993). As shown in Fig. 3C, PR declines as $\alpha$ decreases, although the drop is shallower than that of $D_{\mathrm{KY}}$. This difference is expected: PR is a second-order statistic that is sensitive to how variance is spread across modes, whereas $D_{\mathrm{KY}}$ is a quantity set by local expansion rates of the flow. Consequently, PR can remain relatively high even when only a few directions in phase space are truly unstable, highlighting complementary information provided by these two dimensionality measures.
@@ 5 Discussion
Further, as demonstrated in Fig. 2, heavy-tailed weight distributions make RNNs more robust to changes in gain, a parameter that may correspond biologically to either the width of synaptic weight distributions or to neural gain modulated by neuromodulatory systems (Waterhouse et al., 1988; Shine et al., 2018). Specifically, we observe that networks with heavier-tailed synaptic weights remain near the edge of chaos over a much wider range of gain values than those with Gaussian connectivity, which is commonly assumed in theoretical studies. This property may be especially valuable for biological systems that operate across multiple states (e.g., sleep and waking (Chaudhuri et al., 2019)) or in non-stationary environments. Such robustness could help explain empirical findings that similar neural activity patterns can arise from vastly different underlying circuit parameters in healthy brains (Prinz et al., 2004; Marder, 2011). Meanwhile, as shown in Fig. 3, heavier tails reduce both the Lyapunov dimension and the participation ratio, indicating that the slow manifold supporting long-lasting activity becomes lower-dimensional. Together, these effects imply a tradeoff: heavy-tailed networks are more stable to perturbations but require more neurons to achieve the same computational capacity, such as memory or temporal integration, compared to Gaussian networks.
@@
Notably, a common empirical observation in neuroscience is that neural population activity tends to evolve within a low-dimensional manifold relative to the large number of neurons recorded. This phenomenon has been observed across cortical and subcortical regions, and is often behaviorally meaningful (Nieh et al., 2021; Bondanelli et al., 2021). Theoretical work suggests that low-dimensionality can arise from constraints imposed by circuit connectivity (Mastrogiuseppe and Ostojic, 2018) or task demands (Gao et al., 2017). Our finding that heavier-tailed RNNs yield lower-dimensional attractors biologically aligns with this widespread phenomenon and supports the idea that anatomical connectivity constrains the expressive capacity of population activity.
@@
The observed robustness-dimensionality tradeoff also offers predictions for which tasks heavy-tailed circuits are best suited. Tasks requiring only low-rank dynamics for reliable integration or pattern generation (e.g., binary decision-making (Brunton et al., 2013) or working memory (Panichello and Buschman, 2021)) may benefit from the extended edge-of-chaos regime provided by heavy-tailed weights. In contrast, tasks that rely on high-dimensional dynamics—such as representing multiple independent memories or generating complex trajectories (e.g., virtual reality navigation (Busch et al., 2024))—may require larger networks or connectivity distributions closer to Gaussian. These predictions are directly testable using emerging connectomic (Dorkenwald et al., 2024; The MICrONS Consortium, 2025) and large-scale recording datasets (Bondy et al., 2024; Manley et al., 2024), which can jointly measure synaptic weight distributions and task-related activity dimensionality.
@@
We acknowledge several limitations: our study used untrained rate-based networks with homogeneous units. Including more biologically realistic features such as spiking dynamics (Kim et al., 2019), Dale’s law (Dale, 1935), cell-type diversity (Zeng, 2022; Yao et al., 2023), and synaptic plasticity (Citri and Malenka, 2008) could modify or refine the observed effects. Furthermore, while our results focused on untrained dynamics, a key next step is to study how learning algorithms interact with the broad critical regime and how trained or reservoir computing heavy-tailed networks perform across a range of tasks (Yang et al., 2019; Driscoll et al., 2024). Such a study would help us to understand and predict, based on connectivity alone, what kinds of computations a brain-like circuit is suited to perform—an important goal as we seek to interpret rich new connectomic datasets and understand how synaptic connectivity ties to function (Garner et al., 2024; Seung, 2024).
@@ Appendix A Mathematical analysis of the transition in annealed networks
where $\gamma^{(t)}$ depends on the perturbation at time $t-1$, and $z_{i}^{(t)}$ are i.i.d. $\alpha$-stable variables. This leads to the recursive formula for scalar $\gamma^{(t)}$
@@
where $\gamma^{(1)}$ is deterministically specified by the input perturbation $\varepsilon^{(0)}$. It is known [Kesten, 1973, Statman et al., 2014] that this sum diverges to $\infty$- ∞ almost surely if $\langle\ln\xi\rangle<0$ and diverges to $\infty$ almost surely if $\langle\ln\xi\rangle>0$. Accordingly, the sequence $\left(\gamma^{(t)}\right)_{t=1}^{\infty}$ either converges to 00 or diverges. Therefore, the critical width of the synaptic weight distribution is given by
@@ Appendix B Derivation of the logarithmic decay of $g^{*}(N)$
[…]
@@
$F_{1,\alpha}(s)=1-A_{\alpha}\left(B_{\alpha}-\ln s\right)+O(s^{2})$ (20)
@@ Appendix F Visualizations of multiple realizations of Lyapunov spectrum
To assess the variability across realizations of network connectivity, we visualize the full Lyapunov spectrum from three independent trials (different seeds) for networks with $N=1000$, across both autonomous and noisy stimulus-driven settings. These spectra are computed near the estimated critical gain $g^{*}$ (obtained in Figs. 2 and 5), where the maximum Lyapunov exponent $\lambda_{\max}$ first crosses zero in each condition. The same value of $\langle g^{*}\rangle$, computed by averaging $g^{*}$ of multiple runs, is used across different seeds in these figures. The actual transition point can vary in each realization. Moreover, due to the finite resolution of the grid of $g$ values used in simulations, $g^{*}$ is overestimated in each seed. Thus, in some realizations, the right edge of the histogram may exceed 00, and the average histograms presented in Figs. 3A and 6A can feature some positive Lyapunov exponents. A more precise estimate could be obtained through a finer-grained or binary search over gain values near the transition point. However, this additional numerical precision would unlikely affect our overall conclusions.
@@
Since we only examine the maximum (top-1) Lyapunov exponent in Fig. 2, the number of top exponents computed (denoted k_LE in the codebase) is irrelevant as long as k_LE $>1$. Throughout Fig. 2 and this appendix, we use the default k_LE $=100$. Additionally, we exclude an initial warmup period before accumulating exponents to avoid contamination from transients (Appendix C). In Fig. 2, the network is run for $T=3000$ steps, and Lyapunov exponents are accumulated over the final $K=100$ steps.
@@
[…]
@@
Appendix J Additional details
J.1 Experiments compute resources
@@
J.2 Licenses for existing assets

### v2

@@ 4.3.1 Lyapunov spectrum shows compressed slow manifold in heavy-tailed RNNs
Second, we evaluate the participation ratio (PR), a widely used metric for gauging the effective dimensionality of neural population activity. It has been leveraged to quantify task-relevant low-dimensional subspaces and other properties of multi-unit neuronal recordings in behaving animals (Gao et al., 2017), and summarize the collective modes visited by recurrent spiking networks and reveal how these modes depend on local connectivity motifs (Recanatesi et al., 2019). PR measures how variance in population activity is distributed across the eigenmodes of the covariance matrix, providing a compact read‑out of the number of degrees of freedom the network explores (Kramer and MacKinnon, 1993). As shown in Fig. 3C, PR declines as $\alpha$ decreases, although the drop is shallower than that of $D_{\mathrm{KY}}$. This difference is expected: PR is a second‑order statistic that is sensitive to how variance is spread across modes, whereas $D_{\mathrm{KY}}$ is a quantity set by local expansion rates of the flow. Consequently, PR can remain relatively high even when only a few directions in phase space are truly unstable, highlighting complementary information provided by these two dimensionality measures.
We hypothesize that the large disparity in Lyapunov dimensions between Gaussian and heavy-tailed networks arises from the broader dispersion of Lyapunov exponents in the latter as shown in Fig. 3A. Intuitively, only a small subset of leading exponents becomes positive near the edge of chaos in heavy-tailed networks, resulting in a lower overall Lyapunov dimension. This effect likely reflects the more heterogeneous eigenvalue distribution of the underlying weight matrix. However, the precise mapping between the weight matrix spectrum and the Jacobian’s Lyapunov spectrum remains nontrivial and warrants further analysis.
@@ 5 Discussion
Further, as demonstrated in Fig. 2, heavy‑tailed weight distributions make RNNs more robust to changes in gain, a parameter that may correspond biologically to either the width of synaptic weight distributions or to neural gain modulated by neuromodulatory systems (Waterhouse et al., 1988; Shine et al., 2018). Specifically, we observe that networks with heavier-tailed synaptic weights remain near the edge of chaos over a much wider range of gain values than those with Gaussian connectivity, which is commonly assumed in theoretical studies. This property may be especially valuable for biological systems that operate across multiple states (e.g., sleep and waking (Chaudhuri et al., 2019)) or in non-stationary environments. Such robustness could help explain empirical findings that similar neural activity patterns can arise from vastly different underlying circuit parameters in healthy brains (Prinz et al., 2004; Marder, 2011). Meanwhile, as shown in Fig. 3, heavier tails reduce both the Lyapunov dimension and the participation ratio, indicating that the slow manifold supporting long-lasting activity becomes lower-dimensional. Our further analysis show that a handful of “mega-synapses” drives the dynamics, implying the robustness and low-dimensionality largely stem from extreme outliers (Appendix K). Together, these effects imply a tradeoff: heavy-tailed networks are more stable to perturbations but require more neurons to achieve the same computational capacity, such as memory or temporal integration, compared to Gaussian networks.
@@
Notably, a common empirical observation in neuroscience is that neural population activity tends to evolve within a low-dimensional manifold relative to the large number of neurons recorded. This phenomenon has been observed across cortical and subcortical regions, and is often behaviorally meaningful (Nieh et al., 2021; Bondanelli et al., 2021). Theoretical work suggests that low-dimensionality can arise from constraints imposed by circuit connectivity (Mastrogiuseppe and Ostojic, 2018) or task demands (Gao et al., 2017). Our finding that heavier-tailed RNNs yield lower-dimensional attractors biologically aligns with this widespread phenomenon and provides evidence that anatomical connectivity might constrain the expressive capacity of population activity.
@@
The observed robustness-dimensionality tradeoff also offers predictions for which tasks heavy-tailed circuits can be best suited. Tasks requiring only low-rank dynamics for reliable integration or pattern generation (e.g., binary decision-making (Brunton et al., 2013) or working memory (Panichello and Buschman, 2021)) may benefit from the extended edge-of-chaos regime provided by heavy-tailed weights. In contrast, tasks that rely on high-dimensional dynamics—such as representing multiple independent memories or generating complex trajectories (e.g., virtual reality navigation (Busch et al., 2024))—may require larger networks or connectivity distributions closer to Gaussian. These predictions can be tested using emerging connectomic (Dorkenwald et al., 2024; The MICrONS Consortium, 2025) and large-scale recording datasets (Bondy et al., 2024; Manley et al., 2024), which can jointly measure synaptic weight distributions and task-related activity dimensionality.
Our framework can be naturally extended to a mixture setting, improving biological plausibility. For instance, neurons could be homogeneous and each draw weights randomly from one of multiple heavy-tailed distributions or form interacting subpopulations with distinct $\alpha$ values and connectivity motifs. Such extensions may capture diversity across neuronal cell types (Jin et al., 2025; Zeng, 2022) and offer a promising direction for future work.
@@
We acknowledge several limitations: our study used untrained rate-based networks with homogeneous units. Including more biologically realistic features such as spiking dynamics (Kim et al., 2019), Dale’s law (Dale, 1935), cell-type diversity (Zeng, 2022; Yao et al., 2023), and synaptic plasticity (Citri and Malenka, 2008) could modify or refine the observed effects. Furthermore, while our results focused on untrained dynamics, a key next step is to study how learning algorithms interact with the broad critical regime and how trained or reservoir computing heavy-tailed networks perform across a range of tasks (Yang et al., 2019; Driscoll et al., 2024). Such a study would help us to understand and predict, based on connectivity alone, what kinds of computations a brain-like circuit is suited to perform—an important goal as we seek to interpret rich new connectomic datasets and understand how synaptic connectivity ties to function (Garner et al., 2024; Seung, 2024). Another valuable next step is to extend this work toward direct comparison with neural recordings. For example, future studies could estimate Lyapunov spectra or related dynamical signatures from long, high-resolution neural activity trajectories. While such analyses are technically challenging and require stable, extended recordings, they would offer a powerful bridge between theory and experiment.
@@
6 Acknowledgements
Eva Yi Xie was supported by a NeurIPS Foundation travel award for the NeurIPS conference presentation. Support for Stefan Mihalas was provided in part by NSF (2223725) and NIH (RF1DA055669); support for Łukasz Kuśmierz was provided in part by NSF (2223725). We are grateful to the colleagues at the Allen Institute and the COSYNE 2025 community for feedback on an early version of this work.
@@ Appendix A Mathematical analysis of the transition in annealed networks
where $\gamma^{(t)}$ depends on the perturbation at time t1t-1, and $z_{i}^{(t)}$ are i.i.d. $\alpha$-stable variables. This leads to the recursive formula for scalar $\gamma^{(t)}$
@@
where $\gamma^{(1)}$ is deterministically specified by the input perturbation $\varepsilon^{(0)}$. It is known [Kesten, 1973, Statman et al., 2014] that this sum diverges to -$\infty$ almost surely if $\langle\ln\xi\rangle<0$ and diverges to $\infty$ almost surely if $\langle\ln\xi\rangle>0$. Accordingly, the sequence $\left(\gamma^{(t)}\right)_{t=1}^{\infty}$ either converges to 0 or diverges. Therefore, the critical width of the synaptic weight distribution is given by
@@ Appendix B Derivation of the logarithmic decay of $g^{*}(N)$
$\left\langle e^{-s|x|^{\alpha}}\right\rangle_{z\sim\rho_{z}}=s\int\limits_{s}^{\infty}duu^{-2}e^{-u}=s\Gamma(-1,s)\approx 1-s\left(1-\gamma-\ln s\right)+O(s^{2})$ (19)
@@
$F_{1,\alpha}(s)=1-A_{\alpha}s\left(B_{\alpha}-\ln s\right)+O(s^{2})$ (20)
@@ Appendix F Visualizations of multiple realizations of Lyapunov spectrum
To assess the variability across realizations of network connectivity, we visualize the full Lyapunov spectrum from three independent trials (different seeds) for networks with $N=1000$, across both autonomous and noisy stimulus-driven settings. These spectra are computed near the estimated critical gain $g^{*}$ (obtained in Figs. 2 and 5), where the maximum Lyapunov exponent $\lambda_{\max}$ first crosses zero in each condition. The same value of $\langle g^{*}\rangle$, computed by averaging $g^{*}$ of multiple runs, is used across different seeds in these figures. The actual transition point can vary in each realization. Moreover, due to the finite resolution of the grid of $g$ values used in simulations, $g^{*}$ is overestimated in each seed. Thus, in some realizations, the right edge of the histogram may exceed 0, and the average histograms presented in Figs. 3A and 6A can feature some positive Lyapunov exponents. A more precise estimate could be obtained through a finer-grained or binary search over gain values near the transition point. However, this additional numerical precision would unlikely affect our overall conclusions.
@@
Since we only examine the maximum (top-1) Lyapunov exponent in Fig. 2, the number of top exponents computed (denoted k_LE in the codebase) is irrelevant as long as k_LE >1>1. Throughout Fig. 2 and this appendix, we use the default k_LE =100=100. Additionally, we exclude an initial warmup period before accumulating exponents to avoid contamination from transients (Appendix C). In Fig. 2, the network is run for $T=3000$ steps, and Lyapunov exponents are accumulated over the final $K=100$ steps.
@@
We use the full Lyapunov spectrum to compute the results shown in Fig. 3B due to the definition of Lyapunov dimension (Eqn. 4), hence $\texttt{k\_LE}=N=1000$ in Fig. 3B. We simulate the dynamics over a total number of $T=2950$ steps, and use the last $K=50$ steps to compute the Lyapunov dimension.
@@
Appendix J Behavior of the quenched transition point as a function of NN
In finite-sized quenched networks, the location of the transition point fluctuates between realizations of the weight matrix. Since our annealed theory does not offer any insight into the nature of these fluctuations, we resorted to numerical simulations to study how the statistics of $g^{*}$ scale with $N$. The results for the representative case of $\alpha=1$ are shown in Fig. 17. The mean location of the transition point scales like $1/\ln N$, in line with our theoretical predictions (Fig. 17A). The annealed prediction seems to underestimate the true mean over quenched realizations. The standard deviation of $g^{*}$ decreases with $N$ at a comparable rate as the mean (Fig. 17B). The coefficient of variation of $g^{*}$ falls off slowly in the studied range of $N$ (Fig. 17C), suggesting that the location of the transition may be (weakly) self-averaging [Wiseman and Domany, 1998] in this system.
[…]
Figure 17: Statistics of $g^{*}$ in quenched networks as functions of $N$. Note the logarithmic scale on the x-axis. Each point corresponds to the statistics estimated using $100$ independent realizations of the weight matrix. For each value of $N$, we included three data points that correspond to independent estimates calculated based on different random seeds. (A) Reciprocal of the mean. (B) Standard deviation. (C) Coefficient of variation.
Appendix K The effect of mega-synapses on dynamics
To test whether robustness and low dimensionality arise from global heavy-tailed statistics or a few extreme “mega-synapses,” we pruned recurrent weights in a network of size $N=1000$ by absolute magnitude (bottom 95%, top 1%, top 3%) and report the results averaged across three trials. The slow transition vanished when top outliers were removed, shifting the critical gain $g^{*}$ to larger values, whereas pruning the weakest 95% had little effect (Fig. 18).
Similarly, removing the bottom percentage of weights has very little effect on the general trend of attractor dimensionality. However, when top outliers are removed, the changes are more nuanced: the attractor dimensionality for heavy-tailed weights increases in the chaotic regime, while the transition to chaos is pushed to a larger $g^{*}$ when more top outlier weights are pruned as mentioned above (Figs. 1920). The general ranking of dimensionality by $\alpha$ is largely consistent with the main paper for both dimension measures, though the max dimensionality of $\alpha=1.5$ is comparable to that of $\alpha=2$ over a range of $g$ when top outliers are pruned.
[…]
Figure 18: Effect of pruning on critical gain $g^{*}$. (A) bottom 95% removed. (B) top 1% removed. (C) top 3% removed.
[…]
Figure 19: Effect of pruning on Lyapunov dimension. (A) bottom 95% removed. (B) top 1% removed. (C) top 3% (LD) removed.
[…]
Figure 20: Effect of pruning on participation ratio (PR). (A) bottom 95% removed. (B) top 1% removed. (C) top 3% removed.
Appendix L Information processing in heavy-tailed recurrent neural networks
To examine whether our results extend to structured external inputs (and toward learned settings), we provide a proof-of-concept through a reservoir-computing experiment on the delayed-memory XOR task (a similar task is used in [Huh and Sejnowski, 2018]). We use networks of size $N=1000$ and report average performance across three trials. Specifically, in the XOR task, each trial presents two binary stimulus vectors $s_{1},s_{2}$ separated by silent delays; after the second delay, the readout must report $\mathrm{XOR}(s_{1},s_{2})$, requiring short-term maintenance of both stimuli and a nonlinear decision rule.
Across gains $g$, heavy-tailed reservoirs exhibited a broader and more stable operating regime than Gaussian reservoirs: the transition to chaos was slower and more robust (Fig. 21A), and task performance remained high over a wider range of $g$ (Fig. 21B). These observations suggest that the extended critical regime of heavy-tailed networks can enhance robustness and performance without fine-tuning, with potential benefits for machine learning applications.
[…]
Figure 21: Delayed-memory XOR with heavy-tailed reservoirs. (A) Dynamics across gain $g$ as measured by maximum Lyapunov exponent for Gaussian (pink) vs. heavy-tailed reservoirs (blue), showing a slower, more robust transition to chaos in the latter. (B) Task accuracy of a linear readout on the same reservoirs, with heavy-tailed networks maintaining strong performance over a broader range of $g$.
Appendix M Additional details
M.1 Experiments compute resources
@@
M.2 Licenses for existing assets